OpenStudy (anonymous):

You only have $1500 to invest and the highest interest rate you can find is 3% compounded annually. If you decide to wait 4 years to go on a trip, how much money will you have to spend on a trip? Use the compound interest formula A = P (1 + i)n.

Parth (parthkohli):

You'll have to use \(P( 1+ r)^n\)

Parth (parthkohli):

Here it'd be something like: \( \color{Black}{\Rightarrow 1500(1 + 0.03)^4 }\)

Parth (parthkohli):

@thejoshfish now a calculator haha

Parth (parthkohli):

Just type this into the calculator. \( \color{Black}{\Rightarrow 1500 \times 1.03 \times 1.03 \times 1.03 \times 1.03 }\)

OpenStudy (anonymous):

and the answer is 1688.263215! thanks loads c:

Parth (parthkohli):

:)